I don't know how to solve the following and then graph the answer...
y = 2x - 6
x + y = 0

Answer by mananth(16946)

-2 x + 1 y = -6 .............1
1 x + 1 y = 0 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
2 x -1 y = 6
1 x 1 y = 0
Add the two equations
3 x = 6
/ 3
x = 2
plug value of x in (1)
-2 x + 1 y = -6
-4 + 1 y = -6
1 y = -6 + 4
1 y = -2
y = -2
equation1
when x= 0 y= -6 ( 0 , -6 )
when y = 0 x= 3 ( 3 , 0 )
Equation 2
when x= 0 y= 0 ( 0 , 0 )
when y = 0 x= 0 ( 0 , 0 )

m.ananth@hotmail.ca